OpenStudy (anonymous):

A library expansion, begun in 2002, was expected to cost $107 million. By 2006, library officials estimate the cost would be $32 million over budget. By what percent did the actual cost exceed the initial estimate?

OpenStudy (anonymous):

I need major help!! I have tried and tried to understand this problem and I just cant seem to get it!!

OpenStudy (anonymous):

Maybe easier numbers would help you get the idea... Say you buy a meal for $10. Next time, you buy a meal for $15. By what percent did your second meal price increase over your first meal price? It increased by $5, right? You compare the new price $15 to the old price $10, and subtract to see that the price increased by $5. Then divide that price difference by the original price to get the percentage increase... 5/10 = 50% increase. Percent increase = (new amount - old amount) / (old amount)

OpenStudy (anonymous):

ok, I see it.. So since the question is saying the budget went over would I add the 32 million to the 107 million to get 139 million instead of subtracting? Or am I wrong?

OpenStudy (anonymous):

well, you confused yourself just a bit, but you're on the right track :) new amount = 32 + 107 = 139 since the problem said it would be over budget by $32 million. However, when you do "new - old", you get 139 - 107 = 32 So the percentage change is (139 - 107)/107 you could have just used 32/107 and gotten the same result.

OpenStudy (anonymous):

Oh I see, so where I was confusing myself was just the subtraction.. Ok, I think I get it you explained it very well. Thank you very much for your help!

OpenStudy (anonymous):

Now i just had to round my answer to the nearest tenth and I did get the correct answer finally! Thanks so much!

OpenStudy (anonymous):

great! glad to help :)